# From Memorization to Creativity: LLM as a Designer of Novel Neural-Architectures

### v1

@@ Abstract
Large language models (LLMs) excel in program synthesis, yet their ability to autonomously navigate neural architecture design—balancing syntactic reliability, performance, and structural novelty—remains underexplored. We address this by placing a code-oriented LLM within a closed-loop synthesis framework, analyzing its evolution over 22 supervised fine-tuning cycles. The model synthesizes PyTorch convolutional networks which are validated, evaluated via low-fidelity performance signals (single-epoch accuracy), and filtered using a MinHash–Jaccard criterion to prevent structural redundancy. High-performing, novel architectures are converted into prompt–code pairs for iterative fine-tuning via parameter-efficient LoRA adaptation, initialized from the LEMUR dataset. Across cycles, the LLM internalizes empirical architectural priors, becoming a robust generator. The valid generation rate stabilizes at 50.6% (peaking at 74.5%), while mean first-epoch accuracy rises from 28.06% to 50.99%, and the fraction of candidates exceeding 40% accuracy grows from 2.04% to 96.81%. Analyses confirm the model moves beyond replicating existing motifs, synthesizing 455 high-performing architectures absent from the original corpus. By grounding code synthesis in execution feedback, this work provides a scalable blueprint for transforming stochastic generators into autonomous, performance-driven neural designers, establishing that LLMs can internalize empirical, non-textual rewards to transcend their training data.
From Memorization to Creativity:
LLM as a Designer of Novel Neural-Architectures
Waleed Khalid   Dmitry Ignatov   Radu Timofte Computer Vision Lab, CAIDAS & IFI, University of Würzburg, Germany
@@ 1 Introduction
Designing effective neural network architectures remains a central bottleneck in modern deep learning. Neural Architecture Search (NAS) emerged to automate this process through reinforcement learning, evolutionary algorithms, and differentiable optimization (zoph2017neural; white2023neural; kang2023neural). While successful, traditional NAS often incurs prohibitive computational costs. Consequently, the CIFAR-10 classification task (krizhevsky2009learning) has become a canonical benchmark for evaluating these automated design strategies.
@@
In parallel, large language models (LLMs) have revolutionized program synthesis, enabling the generation of complex source code from natural-language instructions. Recent frameworks like LLMatic and LEMONADE have begun leveraging LLMs to emit full network definitions, demonstrating their potential as architecture generators (nasir2024llmatic; rahman2025lemonade). However, existing studies primarily focus on final model accuracy and search efficiency, offering limited insight into the generator’s reliability. Specifically, it remains unclear how LLM-driven synthesis evolves under iterative refinement, particularly regarding syntactic validity, structural novelty, and the ability to maintain diversity as the model specializes.
@@
In this work, we therefore consider an LLM purely as an architecture synthesizer and address the following central question: if we repeatedly fine-tune an LLM on its own successful generations, does its ability to produce valid, high-quality, and structurally novel network architectures measurably improve over time? Rather than optimizing for final test accuracy after long training runs, we deliberately adopt a low-cost performance proxy: the classification accuracy achieved after a single training epoch on CIFAR-10 (krizhevsky2009learning). This early-epoch accuracy is inexpensive to obtain and directly reflects how well the generated architectures support fast initial learning. At the same time, we treat structural uniqueness as a first-class objective, because practical NAS workflows benefit not only from strong individual models but also from diverse candidates that explore different regions of the design space (white2023neural; kang2023neural).
@@
Concretely, we execute an LLM-driven synthesis loop over 22 cycles where candidate architectures are filtered for compilation validity, trained for a single epoch, and subjected to MinHash–Jaccard novelty analysis. Our results demonstrate that this iterative generate–evaluate–select–fine-tune process, guided by low-fidelity signals, monotonically improves both generator reliability and model performance while maintaining significant structural diversity, effectively reshaping the LLM into a robust architectural prior.
In summary, This work advances the intersection of automated program synthesis and neural architecture design through three primary contributions. First, we establish an LLM-driven synthesis framework that treats the generator as a trainable architectural prior, optimizing for a triad of objectives: syntactic validity, early-epoch performance, and structural novelty. Second, we introduce a code-level novelty filter utilizing MinHash–Jaccard similarity to programmatically ensure meaningful design-space expansion. Third, we provide a 22-cycle longitudinal analysis demonstrating that iterative fine-tuning significantly enhances generation reliability and model quality without sacrificing architectural diversity.
@@ 2 Related Work
The development of Neural Architecture Search (NAS) has significantly automated network design through reinforcement learning, evolutionary algorithms, and differentiable optimization, though often at a prohibitive computational cost (elsken2019nas; white2023neural; kang2023neural). To ameliorate the expense of repeated candidate training, the field has increasingly relied on low-fidelity proxies, including early-stopped training, learning-curve extrapolation, and training-free zero-cost signals (domhan2015lcextrapation; ru2020speedy; zela2020darts). While our work utilizes single-epoch accuracy as an efficient performance proxy, we diverge from traditional NAS by employing these signals to shape the behavioral priors of a generative Large Language Model (LLM) rather than optimizing within a static, handcrafted search space.
@@
In parallel, the advent of code-capable LLMs has introduced a paradigm shift toward synthesizing complete model implementations from natural language. Frameworks such as LLMatic have demonstrated the efficacy of coupling LLM-driven mutation with quality-diversity search (nasir2024llmatic), while others have integrated iterative refinement to satisfy stringent deployment constraints (rahman2025lemonade). More recent self-improving systems, such as SEKI and RZ-NAS, leverage performance-guided evolution or reflective reasoning to improve design outcomes (cai2025seki; ji2025rznas). Despite these advances, existing research typically evaluates success through final search efficiency or peak accuracy. Our approach provides a distinct longitudinal perspective by explicitly characterizing the evolution of the generator itself—tracking metrics of validity, performance distribution shifts, and code-level novelty across twenty-two successive cycles of supervised fine-tuning.
@@
Crucially, the utility of LLM-generated architectures is predicated on both functional reliability and structural diversity. While standard code generation benchmarks prioritize functional correctness and unit-test pass rates (chen2021codexeval; jiang2024codellmsurvey), the structured nature of PyTorch programs necessitates a more nuanced separation between executable validity—comprising parsing, instantiation, and forward passes—and downstream learning quality. To prevent the collapse of the generator into redundant motifs or trivial rewrites, we incorporate a MinHash–Jaccard near-duplicate filter. This ensures that the fine-tuning corpus is augmented only with implementations that are both performant and structurally novel, fostering a diverse architectural prior that transcends simple memorization of existing training samples.
@@ 3 Method
We treat a code-oriented large language model (LLM) as a stochastic generator of neural network architectures and study how its behavior changes under an iterative refinement loop. We run $22$ synthesis cycles indexed by $c\in\{1,\dots,22\}$. In each cycle, the LLM generates candidate PyTorch models; we execute validity checks, run a fixed first-epoch CIFAR-10 (krizhevsky2009learning) training protocol to obtain a low-cost performance proxy, filter for novelty, and then fine-tune the LLM on the accepted outputs before proceeding to the next cycle. Figure 1 summarizes the generate–evaluate–select–fine-tune loop.
@@
A code-focused LLM is used as a conditional sampler over PyTorch architectures. Prompts specify CIFAR-10 classification, input/output shapes (e.g., $(N,3,32,32)\rightarrow 10$ logits), and constraints on permissible operations (standard conv/pool/norm/activations; no pretrained weights or external feature extractors). A parameter budget of at most 500,000500{,}000 parameters is imposed alongside an implicit edge-friendly latency target. Each generated model is required to implement a fixed API contract: a class Net(nn.Module) with __init__, forward, train_setup, learn, and a function supported_hyperparameters() returning {"lr", "momentum"}. Prompts require import torch and import torch.nn as nn and instruct the model to output a single nn.Module definition (no data loading or training loops). The prompt template, decoding configuration, and maximum generation length are held fixed across all cycles to avoid prompt drift; consequently, changes in outputs are attributable to training data and fine-tuning.
Drawing upon recent advances in LLM applications across diverse domains (Gado2025llm; Rupani2025llm; ABrain.NN-RAG) and prior architectural synthesis studies within the NNGPT framework (ABrain.HPGPT; ABrain.NN-Captioning_2025; ABrain.Prompt; ABrain.NNGPT-Fractal; ABrain.Transform), we initialize supervision using the LEMUR Neural Network Dataset as part of the NNGPT ecosystem (ABrain.NNGPT). This dataset constitutes the foundational knowledge base of NNGPT and encompasses a broad range of both high-capacity and edge-optimized neural network models (ABrain.NN-Dataset; ABrain.LEMUR2; ABrain.NN-Lite). Leveraging this diverse collection of performance-annotated metadata enables the model to internalize empirical performance signals and to transition from stochastic code generation toward informed, task-aligned architectural design. A total of 109,913 model records are obtained and then pruned by removing text-level near-duplicates using MinHash/LSH over token shingles. Specifically, each code snippet is tokenized, $k$-shingles are formed over tokens, and MinHash signatures with locality-sensitive hashing (LSH) are used to retrieve near-neighbor candidates efficiently; For every successfully parsed and trained candidate, we compute two Jaccard similarities over token-level shingles of the model code. The first, denoted $J_{\text{train}}$, measures similarity to the deduplicated supervised training set (LEMUR-derived models plus all previously accepted generations); the second, denoted $J_{\text{gen}}^{(c)}$, measures similarity to the set of all code samples generated earlier in the same cycle $c$.
In the implementation, these quantities are logged as nn_jaccard_train and nn_jaccard_gen, respectively, and are accompanied by Boolean flags near_dup_text_train and near_dup_text_gen that indicate whether either similarity exceeds a predefined near-duplicate threshold. A typical log entry for a valid model therefore contains fields such as
"nn_jaccard_train": 0.0,
"near_dup_text_train": false,
"nn_jaccard_gen": 0.8047,
"near_dup_text_gen": false,
"rejection_count": 1
where rejection_count records how many previously sampled candidates were rejected for being too similar to existing code before accepting the current model.
Formally, let $S_{\text{train}}$ denote the collection of shingle sets corresponding to all models in the current supervised corpus, and let $S_{\text{gen}}^{(c)}$ denote the collection of shingle sets for all models generated in cycle $c$ prior to the current candidate. For a new candidate with shingle set $A$, we estimate
@@
via their MinHash signatures and LSH index. If either similarity exceeds a near-duplicate threshold $\tau$ (e.g., $\tau\approx 0.9$), the candidate is marked as a text-level near-duplicate and rejected. Sampling continues until a valid architecture is found that is sufficiently dissimilar to both the current training set and earlier generations in the same cycle.
This procedure yields 1,065 unique records. These records are converted to instruction-following chat examples via a ChatPrepConfig; 216 records are dropped due to conversion/format issues, leaving 849 converted records. For each converted record, two training examples are created with identical assistant code but different user descriptions (MNIST vs. CIFAR-10), yielding $849\times 2=1{,}698$ chat-style prompt–code pairs for the initial training split.
During the 22-cycle loop, the corpus is augmented with self-generated models. At the end of each cycle, candidate models are considered for inclusion if they (i) compile and train, (ii) exceed a first-epoch CIFAR-10 accuracy threshold of 40%, and (iii) pass a near-duplicate filter against both the current training corpus and earlier generations from the same cycle using the same MinHash–Jaccard approach described above. For novelty filtering, each trained candidate with code-shingle set $S$ is assigned an estimated maximum Jaccard similarity to the supervised corpus and to other candidates in the same cycle; near-duplicates above a threshold $\tau\approx 0.9$ are rejected. Each qualifying model is converted into a chat-format prompt–code pair and appended to the training set for the next cycle. Across 22 cycles, this process adds 455 unique high-accuracy models, expanding the training corpus from 1,698 to 2,153 examples by Cycle 22 (Table 1).
Source Prompt–code pairs
LEMUR (deduplicated, chat-format train split) 1,698
Self-generated (22 cycles, $\geq 40\%$ acc., novel) 455
@@
Table 1: Supervised training corpus by the end of cycle 22.
Figure 2: Valid generation rate per cycle (1–22) with Wilson 95% confidence intervals.
Each generated code snippet is executed in an isolated environment. Candidates are rejected if Python parsing fails, if Net cannot be instantiated (e.g., missing arguments or incompatible dimensions), or if a dummy forward pass raises an exception (e.g., tensor shape mismatch). To ensure comparability across experimental cycles, all remaining candidates are subjected to a standardized training protocol on the CIFAR-10 dataset. Key hyperparameters—including the training/validation split, input resolution, optimization schedule, and batch size—are held constant. Advanced data augmentation techniques derived from Aboudeshish2025augmentation are utilized to maintain a consistent baseline. Implementation details for MinHash/LSH near-duplicate detection and novelty filtering (including shingle size, signature length, and thresholds) are provided in Appendix A.
For each syntactically valid model $m$, the top-1 validation accuracy after a single epoch, denoted as $A(m)$, is recorded. This approach serves as a low-fidelity performance proxy, a common practice in NAS to balance computational efficiency with evaluative reliability (ru2020speedy; zela2020understanding). This standardized evaluation allows for the efficient ranking of model candidates while minimizing the computational overhead of full convergence training. Per-cycle summaries include the valid generation rate.
The valid generation rate is
$p_{\mathrm{valid}}^{(c)}=\frac{N_{\mathrm{valid}}^{(c)}}{N_{\mathrm{gen}}^{(c)}}.$ (2)
Let $N_{\mathrm{gen}}^{(c)}$ denote the number of candidate architectures sampled from the LLM in cycle $c$, and let $N_{\mathrm{valid}}^{(c)}$ denote the subset that compile and train successfully.
For all valid models in cycle $c$, let
$\mathcal{A}^{(c)}=\{A(m):m\in\mathcal{M}_{\mathrm{valid}}^{(c)}\}.$
For the sample mean $\bar{A}^{(c)}$ and standard deviation $s^{(c)}$ computed over $n_{c}=|\mathcal{A}^{(c)}|$, we use the usual $t$-based 95% confidence interval:
$\bar{A}^{(c)}\pm t_{0.975,\,n_{c}-1}\frac{s^{(c)}}{\sqrt{n_{c}}}.$ (3)
For any proportion $\hat{p}=k/n$ (e.g., $p_{\mathrm{valid}}^{(c)}$ or the proportion with $A(m)\geq\tau$), we report Wilson score confidence intervals (wilson1927probable).

### v2

# From Memorization to Creativity: LLM as a Designer of Novel Neural Architectures

@@ Abstract
Large language models (LLMs) excel in program synthesis, yet their capacity for neural architecture design—balancing syntactic reliability, performance, and structural novelty—remains underexplored. We present a closed-loop architecture synthesis pipeline within the NNGPT framework, in which a code-oriented LLM evolves over 22 supervised fine-tuning cycles. At each cycle, the LLM synthesizes PyTorch convolutional networks, validated via low-fidelity performance signals and filtered via a MinHash–Jaccard criterion to prevent structural redundancy before being incorporated into the LEMUR dataset. High-performing candidates with novel architectures are converted into prompt–code pairs for parameter-efficient LoRA fine-tuning. This feedback loop drives a measurable distributional shift, progressively internalizing empirical architectural priors such that valid and high-performing outputs evolve from scarce to dominant across cycles. On CIFAR-10, the valid generation rate stabilizes at 50.6% (peaking at 74.5%), mean first-epoch accuracy rises from 28.1% to 51.0%, and candidates exceeding 40% accuracy grow from 2.0% to 96.8%. Cross-dataset transfer to CIFAR-100 and SVHN confirms that improved validity, shifted accuracy distributions, and sustained novelty generalize across benchmarks of varying difficulty and visual domain. Across 22 cycles, 455 unique architectures absent from the original corpus are admitted under the novelty filter. By grounding synthesis in execution feedback and novelty filtering, we demonstrate that iterative self-supervised fine-tuning reshapes an LLM into a task-specialized architectural prior—improving generation reliability, proxy performance, and structural diversity—offering a reproducible, annotation-free alternative to hand-crafted search spaces. All code artifacts are publicly released under https://github.com/ABrain-One, including the code retrieval engine (NN-RAG), prompt generation (NN-Dup), and LLM fine-tuning (NNGPT) pipelines, as well as the generated models (NN Dataset); the fine-tuned LLM is available at https://huggingface.co/ABrain/NNGPT-UniqueArch-Rag.
@@ 1 Introduction
Designing effective neural network architectures remains a central bottleneck in modern deep learning. Neural Architecture Search (NAS) emerged to automate this process through reinforcement learning, evolutionary algorithms, and differentiable optimization [41, 38, 20]. While successful, traditional NAS often incurs prohibitive computational costs. Consequently, the CIFAR-10 classification task [25] has become a canonical benchmark for evaluating these automated design strategies.
@@
In parallel, large language models (LLMs) have revolutionized program synthesis, enabling the generation of complex source code from natural-language instructions. Recent frameworks like LLMatic and LEMONADE have begun leveraging LLMs to emit full network definitions, demonstrating their potential as architecture generators [30, 32]. However, existing studies primarily focus on final model accuracy and search efficiency, offering limited insight into the generator’s reliability. Specifically, it remains unclear how LLM-driven synthesis evolves under iterative refinement, particularly regarding syntactic validity, structural novelty, and the ability to maintain diversity as the model specializes.
@@
In this work, we therefore consider an LLM purely as an architecture synthesizer and address the following central question: if we repeatedly fine-tune an LLM on its own successful generations, does its ability to produce valid, high-quality, and structurally novel network architectures measurably improve over time? Rather than optimizing for final test accuracy after long training runs, we deliberately adopt a low-cost performance proxy: the classification accuracy achieved after a single training epoch on CIFAR-10 [25]. This early-epoch accuracy is inexpensive to obtain and directly reflects how well the generated architectures support fast initial learning. At the same time, we treat structural uniqueness as a first-class objective, because practical NAS workflows benefit not only from strong individual models but also from diverse candidates that explore different regions of the design space [38, 20].
@@
Concretely, we execute an LLM-driven synthesis loop over 22 cycles where candidate architectures are filtered for compilation validity, trained for a single epoch, and subjected to MinHash–Jaccard novelty analysis. Our results demonstrate that this iterative generate–evaluate–select–fine-tune process, guided by low-fidelity signals, produces a pronounced upward shift in the first-epoch accuracy distribution across cycles, while maintaining significant structural diversity as measured by code-level novelty filtering. The valid generation rate, though not monotonic, stabilizes at levels substantially above early-cycle baselines for much of the run, effectively reshaping the LLM into a task-specialized architectural prior.
It is important to delineate the scope of this study: we do not propose a complete NAS method competing on final test accuracy against established search algorithms, nor do we claim that the generated architectures surpass hand-designed baselines after full training. Instead, we study the generator itself—how its output distribution over architectures changes under iterative self-refinement—using first-epoch accuracy as a low-cost, interpretable proxy. The contribution is therefore a characterization of LLM-as-generator dynamics, providing a foundation upon which future work can build by coupling the refined generator with downstream optimization.
In summary, this work advances the intersection of automated program synthesis and neural architecture design through four primary contributions:
  1. 1.
    We establish an LLM-driven synthesis framework that treats the generator as a trainable architectural prior, optimizing for a triad of objectives: syntactic validity, early-epoch performance, and structural novelty.
  2. 2.
    We introduce a code-level novelty filter utilizing MinHash–Jaccard similarity to programmatically ensure meaningful design-space expansion.
  3. 3.
    We provide a 22-cycle longitudinal analysis demonstrating that iterative fine-tuning induces a substantial distributional shift toward higher-performing architectures, with net improvements in generation reliability, without sacrificing architectural diversity.
  4. 4.
    We validate the generality of the observed trends across three benchmarks—CIFAR-10, CIFAR-100, and SVHN—showing that the iterative refinement loop induces consistent distributional shifts regardless of task difficulty, number of classes, or visual domain.
@@ 2 Related Work
The development of NAS has significantly automated network design through reinforcement learning, evolutionary algorithms, and differentiable optimization, though often at a prohibitive computational cost [12, 38, 20]. To ameliorate the expense of repeated candidate training, the field has increasingly relied on low-fidelity proxies, including early-stopped training, learning-curve extrapolation, and training-free zero-cost signals [9, 33, 40]. While our work utilizes single-epoch accuracy as an efficient performance proxy, we diverge from traditional NAS by employing these signals to shape the behavioral priors of a generative LLM rather than optimizing within a static, handcrafted search space.
@@
In parallel, the advent of code-capable LLMs has introduced a paradigm shift toward synthesizing complete model implementations from natural language. Frameworks such as LLMatic have demonstrated the efficacy of coupling LLM-driven mutation with quality-diversity search [30], while others have integrated iterative refinement to satisfy stringent deployment constraints [32]. More recent self-improving systems, such as SEKI and RZ-NAS, leverage performance-guided evolution or reflective reasoning to improve design outcomes [5, 18]. Despite these advances, existing research typically evaluates success through final search efficiency or peak accuracy. Our approach provides a distinct longitudinal perspective by explicitly characterizing the evolution of the generator itself—tracking metrics of validity, performance distribution shifts, and code-level novelty across twenty-two successive cycles of supervised fine-tuning.
@@
Crucially, the utility of LLM-generated architectures is predicated on both functional reliability and structural diversity. While standard code generation benchmarks prioritize functional correctness and unit-test pass rates [6, 19], the structured nature of PyTorch programs necessitates a more nuanced separation between executable validity—comprising parsing, instantiation, and forward passes—and downstream learning quality. To prevent the collapse of the generator into redundant motifs or trivial rewrites, we incorporate a MinHash–Jaccard near-duplicate filter, ensuring that the fine-tuning corpus is augmented only with implementations that are both performant and structurally novel. Table 1 summarizes the key methodological and empirical distinctions between our framework and the most closely related LLM-based architecture generation methods.
Method Valid Rate Proxy Metric Novel Filter Iter. FT Corpus Growth Best Accuracy
LLMatic [30] N/R Final QD arch. No No ${\sim}94\%^{a}$
LEMONADE [32] ${\approx}75$$85\%^{\dagger}$ Constr. None No No N/Ab
SEKI [5] N/R Final Repo. Yes N/R $97.7\%^{a}$
RZ-NAS [18] ${\approx}98\%^{\ddagger}$ Zero-cost None No No N/Rc
Ours 50.6%e 1-ep. MinHash Yes Yesf 64.0%d
Table 1: Comparison of LLM-based architecture generation frameworks (CIFAR-10). N/R = not reported; FT = fine-tuning. a Fully-trained test accuracy (DARTS/NB201 search space). b Optimizes deployment constraints, not peak accuracy. c Uses zero-cost ranking; final accuracy not reported. dFirst-epoch validation accuracy only; not directly comparable to fully-trained results above. e Explicitly tracked across 22 cycles; no other method reports longitudinal validity. f Training corpus grows from 1,698 to 2,153 prompt–code pairs across cycles. Our framework is the only one to jointly track valid generation rate, apply a code-level novelty filter, grow the training corpus iteratively, and fine-tune on low-fidelity proxy feedback—all without manual annotations.
@@ 3 Method
We treat a code-oriented large language model (LLM) as a stochastic generator of neural network architectures and study how its behavior changes under an iterative refinement loop. We run $22$ synthesis cycles indexed by $c\in\{1,\dots,22\}$. In each cycle, the LLM generates candidate PyTorch models; we execute validity checks, run a fixed first-epoch training protocol to obtain a low-cost performance proxy, filter for novelty, and then fine-tune the LLM on the accepted outputs before proceeding to the next cycle. The primary evaluation is conducted on CIFAR-10 [25]; cross-dataset generalization is assessed on CIFAR-100 [25] and SVHN [31] using the same loop with dataset-appropriate accuracy thresholds (Section 3.2). Figure 1 summarizes the generate–evaluate–select–fine-tune loop.
@@
A code-focused LLM is used as a conditional sampler over PyTorch architectures. Prompts specify CIFAR-10 classification, input/output shapes ($(N,3,32,32)\rightarrow 10$ logits), and constraints on permissible operations (standard conv/pool/norm/activations; no pretrained weights). A parameter budget of at most 500,000500{,}000 parameters is imposed. Each generated model must implement a fixed API contract (class Net(nn.Module) with forward, train_setup, learn); the full prompt specification is provided in Suppl. §I. The prompt template, decoding configuration, and maximum generation length are held fixed across all cycles to avoid prompt drift; consequently, changes in outputs are attributable to training data and fine-tuning.
We initialize supervision using the LEMUR Neural Network Dataset [24, 13], a broad collection of performance-annotated neural network implementations spanning high-capacity and edge-optimized designs [37, 8, 26, 21], drawing on prior work in LLM-driven architectural synthesis within the NNGPT ecosystem [22, 23, 17, 10, 29, 34, 36, 14]. After MinHash/LSH deduplication and chat-format conversion (Suppl. §A–C), the corpus yields 849 unique architectures paired with two task descriptions, producing 1,698 prompt–code training examples.
For every successfully parsed and trained candidate, we compute two Jaccard similarities over token-level shingles of the model code. The first, $J_{\text{train}}$, measures similarity to the deduplicated supervised training set; the second, $J_{\text{gen}}^{(c)}$, measures similarity to the set of all code samples generated earlier in the same cycle $c$. Formally, let $S_{\text{train}}$ denote the collection of shingle sets corresponding to all models in the current supervised corpus, and let $S_{\text{gen}}^{(c)}$ denote the collection for all models generated in cycle $c$ prior to the current candidate. For a new candidate with shingle set $A$:
@@
estimated via MinHash signatures and LSH indexing. If either similarity exceeds a near-duplicate threshold $\tau\approx 0.9$, the candidate is rejected. Sampling continues until a valid architecture is found that is sufficiently dissimilar to both the current training set and earlier generations in the same cycle.
During the 22-cycle loop, the corpus is augmented with self-generated models. At the end of each cycle, candidate models are considered for inclusion if they (i) compile and train, (ii) exceed a first-epoch accuracy threshold on the target dataset (40% for CIFAR-10, 20% for CIFAR-100, 70% for SVHN), and (iii) pass the near-duplicate filter. This process adds 455 unique high-accuracy models across 22 cycles, expanding the training corpus from 1,698 to 2,153 examples (Table 2).
Source N models
LEMUR (deduplicated, train split) 1,698
Self-generated ($\geq 40\%$ acc., novel) 455
@@
Table 2: Supervised training corpus by the end of cycle 22 (CIFAR-10 setting).
Each generated code snippet is executed in an isolated environment. Candidates are rejected if Python parsing fails, if Net cannot be instantiated, or if a dummy forward pass raises an exception. All remaining candidates are subjected to a standardized training protocol on the target dataset, with hyperparameters—including the training/validation split, input resolution, optimization schedule, and batch size—held constant. Advanced data augmentation techniques derived from Aboudeshish et al. [2] are utilized to maintain a consistent baseline. Implementation details for MinHash/LSH near-duplicate detection are provided in Suppl. §A.
For each syntactically valid model $m$, the top-1 validation accuracy after a single epoch, $A(m)$, is recorded as the primary performance signal; a comparison with zero-cost proxy alternatives is given in Suppl. §E. Per-cycle summaries include the valid generation rate $p_{\mathrm{valid}}^{(c)}=N_{\mathrm{valid}}^{(c)}/N_{\mathrm{gen}}^{(c)}$, sample mean and standard deviation of first-epoch accuracy, and the proportion of models exceeding a fixed threshold. We report $t$-based 95% confidence intervals for means and Wilson score intervals [39] for proportions; formal definitions are given in Suppl. §K.